OpenStudy (anonymous):

Write the standard form of the equation of the circle that passes through the points (30, -2), (-1, -19) and (-18, 12). Find the center and the radius. I don't know how to do these! Please help :)

OpenStudy (whpalmer4):

Okay, you have 3 points you know to be on the circle, and the standard formula for a circle has 3 constants (2 for the position of the center, 1 for the length of the radius) to be determined. Your lucky day — 3 points will determine the 3 unknown constants :-) \[(x-h)^2 + (y-k)^2 = r^2\]is the standard formula for a circle with radius \(r\) and center at \((h,k)\) Plug each of your 3 known points into that equation, to get an equation in terms of \(x,y,r\). I'll do the first one: \((30,-2)\): \[(30-h)^2 + (-2-k)^2 = r^2\]expanding that, we get\[h^2-60h + k^2 +4k +904= r^2\] Now do that for the other two points. You'll end up with 3 equations in 3 unknowns, and you can solve the system of equations for the 3 unknowns. Each of the equations will have \(r^2\) as the right hand side, so I suggest you use that to help you solve them.

OpenStudy (anonymous):

Ok, makes sense! Thank you!

OpenStudy (whpalmer4):

Happy to check your answer when you have one.